OpenStudy (anonymous):

Two cars start moving from the same point. One travels south at 56 mi/h and the other travels west at 42 mi/h. At what rate is the distance between the cars increasing two hours later?

OpenStudy (jonnyvonny):

Are you in calculus, at the moment?

OpenStudy (anonymous):

yess

OpenStudy (jonnyvonny):

Daang. I vaguely remember how to solve stuff like this. Anyhow, we both know that we need to know the derivative of the speed of "z":

OpenStudy (jonnyvonny):

Just do x^2+y^2=z^2, just the Pythagorean theorem, to solve for z real quick

OpenStudy (jonnyvonny):

Which, you get z, the speed that they're going apart to be as 70. I have the feeling that you're not with me. Are you?

OpenStudy (anonymous):

4900

OpenStudy (jonnyvonny):

Yes, that is 4900 = z^2, get z alone.